Five years ago a girl was one-half as old as her brother was then and now she is three years younger than he is. Find their present ages.
Please explain how you arrived at the answer... Much thanks!!!!!

Answer by scott8148(6628)

let x="girl's present age" and y="brother's present age"

"Five years ago a girl was one-half as old as her brother was then"
__ 2(x-5)=y-5 __ 2x-10=y-5 __ 2x-5=y

"now she is three years younger than he is" __ x=y-3

substituting __ x=(2x-5)-3 __ x=2x-8 __ x+8=2x